What is the volume of 1.5 Tonne of sand of density 2600 kg/m3? (2 marks)



A bead weighs 15 g. You place it in a graduated cylinder that has 20 cm3 of water in it. After placing the bead in the cylinder, the water level is now at 30 cm3. What is the density of the water?



(2 marks)
Calculate the mass of air in a room of floor dimension 18m × 16m and a height of 6m? Density of air is 1.26 kg/m3






(2 marks)


A piece of charcoal has a volume of 15 cm3 and a mass of 27g. What is its density:
In g/cm3 (2 marks)



In kg/m3 (2 marks)
160 cm= mm
2000 mg = __________________ kg
198 mL =____________________L
338 µg = ___________________ mg
5763 dm = ________________dam

A builder has to fix 3 walls that have fallen down. He needs 2kg of cement for each wall. How much will he need altogether?
In kilograms (2 marks)
1
Expert's answer
2020-06-22T11:08:41-0400

1)


"V=\frac{1500}{2600}=0.58\ m^3"

2)


"\rho=\frac{15}{30-20}=1.5\frac{g}{cm^3}"

3)


"m=(1.26)(18)(16)(6)=2177\ kg"

4)


"\rho=\frac{27}{15}=1.8\frac{g}{cm^3}"

"\rho=1.8\frac{g}{cm^3}=1800\frac{kg}{m^3}"

5)


"160\ cm= 1600\ mm\\

2000\ mg =0.002\ kg\\

198\ mL =0.198\ L\\

338\ µg =0.338\ mg\\

5763\ dm =57.63\ dam"

6)


"m=3(2)=6\ kg"